Is the sum or product of two irrational numbers always irrational?


1
Expert's answer
2021-01-19T18:14:30-0500
  1. The sum of two irrational numbers is not necessarily irrational. For example, "\sqrt{2}\text{ and }-\sqrt{2}" are two irrational numbers, but "\sqrt{2}+(-\sqrt{2})=0" is rational.
  2. The product of two irrational numbers is not necessarily irrational. For example "\sqrt{2}.\sqrt{2}=2" is rational.